# CTR-Driven Ad Text Generation via Online Feedback Preference Optimization

### v2

@@ Introduction
Advertising text is crucial in online advertising, appearing in almost every ad. Effective ad texts not only feature well-crafted expression and accurate conveyance of item information (Murakami et al. 2022), but also achieve strong performance in terms of click-through rate (CTR). Large Language Models (LLMs) enable automated ad text generation (ATG) (Mita et al. 2024) from basic item information and simple instructions, offering significantly higher efficiency than manual creation. Some works (Golobokov et al. 2022; Chai et al. 2022; Jin et al. 2023; Kamigaito et al. 2024) focus on offline metrics like diversity or fluency to enhance the quality of generated ad texts. However, these methods do not guarantee higher CTR performance compared to human-crafted texts, revealing a gap between offline metrics optimization and real-world deployment effectiveness.
@@
Figure 1: Challenges in high-CTR ad text generation. (a) Poor diversity and low CTR performance of sampling-based decoding strategy. (b) Noisy online feedback due to uneven page view (PV) distribution across items. (c) Our work propose an innovative two-stage framework: (1) diverse ad text sampling and (2) CTR-driven preference optimization tailored for these challenges.
@@
Recently, studies (Hughes, Chang, and Zhang 2019; Kamigaito et al. 2021; Kanungo et al. 2022; Wei et al. 2022; Murakami et al. 2025) have focused on directly optimizing for online metrics, particularly CTR. CTR has long been regarded as a key metric for evaluating online advertising effectiveness, as clicks serve as the prerequisite for subsequent conversion to other revenue-related metrics. These approaches typically generate multiple ad texts via sampling-based decoding from LLMs for online testing. The generation models are subsequently optimized using online feedback through algorithms such as contrastive learning and reinforcement learning.
@@
However, these approaches often suffer from the low quality of sampling-based decoding and the noise in online feedback, making it difficult to achieve a significant CTR advantage over human-crafted ad texts. We conduct some analytical experiments and reveal the challenges to realize the CTR improvement. Figure 1(a) shows that generating multiple ad texts via sampling-based decoding from LLMs results in poor diversity and low CTR performance. Figure 1(b) illustrates data confidence issues arising from the uneven distribution of item page views (PV) in online feedback. Short-term online testing makes it difficult to accumulate sufficient high-PV data, causing substantial noise in online feedback. These challenges constrain further improvements in CTR optimization of ATG tasks.
@@
In this work, we propose CTOP, a novel CTR-driven ad Text generation method via Online feedback Preference optimization. CTOP employs an innovative two-stage framework consisting of diverse ad text sampling and CTR-driven preference optimization for high-CTR ad text generation, as illustrated in Figure 1(c). In the first stage, we design a diverse ad text sampling strategy to address the low quality of sampling-based decoding. Specifically, for each item, the strategy uses retrieval-augmented generation (RAG) to retrieve multiple chain-of-thought (CoT)-augmented ad text exemplars, each serving as a one-shot prompt to generate a candidate ad text. In the second stage, we conduct CTR-driven preference optimization tailored for noisy online feedback. We select winning candidates that outperform their corresponding human-crafted ad texts in online A/B/n testing and construct preference pairs by treating the former as preferred and the latter as dispreferred. Each pair is assigned a training weight calculated from gain and confidence coefficients based on its CTR improvement and reliability.
@@
Extensive experimental results on offline and online metrics demonstrate the effectiveness of CTOP. Our method enables efficient generation of diverse ad texts and better aligns with real-world user feedback during preference optimization. Notably, we have successfully applied our method to the item title generation task on a large-scale online shopping platform across various advertising display scenarios. Since late 2024, our approach achieved a relative increase of 1.11% in click-through rate (CTR) and 1.02% in revenue per mille (RPM) compared to the human-crafted baseline, showcasing its strong applicability and effectiveness in advertising systems. The main contributions of our work can be summarized as follows:
@@
  • We propose a novel ad text generation method which optimizes for CTR via preference optimization from online feedback, closing the gap between between generation quality and online performance of ad texts.
@@
  • We design an innovative two-stage framework, consisting of (1) diverse ad text sampling via one-shot in-context learning, using RAG to provide exemplars with CoT reasoning; and (2) CTR-driven preference optimization method that dynamically weighs training pairs by gain and confidence coefficients based on online feedback.
@@
However, evaluating whether a generated ad text achieves “high CTR” is challenging without a meaningful baseline. In practice, each item has a human-crafted ad text $y_{\text{hu}}$ typically written by merchants or the advertising platforms based on domain expertise or empirical knowledge. These human-crafted ad texts generally achieve a certain level of quality and CTR, which serves as a strong and interpretable baseline. Therefore, any generated ad text $y$ that outperforms $y_{\text{hu}}$ in online testing can be considered achieving high CTR. To this end, we reformulate the objective as maximizing the CTR improvement over human-crafted ad texts:
@@
Optimizing ad text generation for high CTR aligns naturally with reinforcement learning, where the goal is to maximize user engagement in real-world environments. In our setting, online A/B/n testing offers an off-policy dataset comprising both LLM-generated and human-crafted texts. Moreover, CTR feedback is delayed, noisy, and requires significant online accumulation, making it challenging to train an explicit reward model for CTR prediction or preference estimation. These factors motivate us to adopt off-policy RL methods with implicit reward modeling (Rafailov et al. 2023; Pal et al. 2024; Wu et al. 2024; Azar et al. 2024; Kim et al. 2024; Ethayarajh et al. 2024; Liu et al. 2025).
Here we adopt DPO (Rafailov et al. 2023) due to its simplicity and efficiency. DPO operates on preference tuples $(x,y_{w},y_{l})$, where $y_{w}$ and $y_{l}$ are the preferred and dispreferred responses with input context $x$. Given a policy model $\pi_{\theta}$ and a reference model $\pi_{\mathrm{ref}}$, the DPO objective is:
@@
We construct preference pairs with each winning candidate as $y_{w}$ and its human-crafted ad text as $y_{l}$. A Qwen-2.5-7B model is supervised fine-tuned on diverse ad texts as $\pi_{\mathrm{ref}}$. With a dataset of 100,000 items from online A/B/n testing, we train a DPO-optimized model. However, A/B testing against human-crafted ad texts yields only a 41.2% win rate.
@@
In this section, we introduce CTOP, a novel CTR-driven ad Text generation method via Online feedback Preference optimization. To ensure meaningful improvements in online performance, we formulate the task objective as maximizing the CTR improvement over human-crafted ad texts, defined in Equation 2. As illustrated in Figure 4, our method employs an innovative two-stage framework of diverse ad text sampling and CTR-driven preference optimization tailored for the two challenges. To address the low-quality issue of sampling-based decoding, we design a diverse ad text sampling strategy in the first stage. To better utilize noisy online feedback, we perform CTR-driven preference optimization in the second stage. The following subsections elaborate on each stage of our proposed method.
@@
As depicted in Challenge I, candidate ad text generation via sampling-based decoding causes poor diversity and inferior CTR performance. This motivates the design of a strategy that achieves both high quality and true diversity.
@@
However, direct imitation of retrieved ad text often fails to capture their stylistic essence. We prompt the model to first analyze the retrieved exemplar and extract imitation tips before ad text generation. This chain-of-thought (CoT) reasoning process yields improvements in the quality of the generated ad texts. To address the inefficiency of online CoT during inference, we pre-generate offline CoTs for all exemplars in the repository. By retrieving exemplars with their CoTs, the in-context learning maintains quality while speeding up ad text generation.
Owing to the inherent stylistic diversity among high-quality exemplars in the repository, our strategy is able to maintain both quality and diversity even as the number of candidates per item $K$ increases. Moreover, this diversity is controllable: our strategy can generate new types of ad texts simply by collecting additional exemplars without extra training. Compared to methods that demand manual prompt engineering for each new style, our approach offers a more scalable and efficient solution with minimal overhead.
@@
Following our sampling strategy, we deploy the generated candidates and the human-crafted ad texts for short-term online A/B/n testing, as shown in Figure 4(b). A candidate that achieves higher CTR than its corresponding human-crafted ad text is defined as a winning candidate. We construct preference pairs with each winning candidate as $y_{w}$ and its human-crafted ad text as $y_{l}$. However, our analysis in Challenge II reveals that noisy online feedback negatively impacts the quality of preference optimization. To address this issue, we propose a CTR-driven preference optimization approach illustrated in Figure 4(c), where each pair is assigned a dynamic weight calculated from gain and confidence coefficients.
@@
We conduct online experiments on a large-scale e-shopping platform, with the task of generating item titles across display advertising scenarios. For diverse ad text sampling, we first build a high-quality item title repository with offline CoTs. Then we retrieve the top-5 semantically relevant exemplars per item and use each as a one-shot prompt of style imitation. A pre-trained Qwen-2.5-7B is used to generate 5 candidate titles for the current item through in-context learning. We select 200,000 items across different categories for diverse generation, resulting in a total of 1 million ad titles. These candidates, together with the human-crafted item titles, are deployed for a two-week online A/B/n testing using 5% of platform traffic as the experimental group. A parallel AA group is simultaneously deployed with identical content and traffic allocation. Due to business sensitivity concerns, we are unable to publicly release the online testing data used in this study.
For CTR-driven preference optimization, we first construct a preference dataset of 500K pairs by pairing each winning candidate and its corresponding human-crafted title from the online A/B/n testing. Each pair is assigned a dynamic weight based on its gain coefficient (Equation 5, with scaling factor $m=2.5$) and confidence coefficient (Equation 6). We use a Qwen-2.5-7B model supervised fine-tuned on diverse titles in the A/B/n testing as the reference model $\pi_{\text{ref}}$, which also initializes the policy model $\pi_{\theta}$. We optimize the objective in Equation 7 for one training epoch with a learning rate of $2\times 10^{-5}$. All training processes are conducted on machines equipped with 8 NVIDIA A100 GPUs.
@@
  • Top-1 sampling: Our diverse ad text sampling with sampled number $K=1$.
@@
training-free
Prompt-only 32.1 -9.83
@@
For training-free baselines (Prompt-only and Top-1 sampling), both underperform the human-crafted ad texts. Prompt-only yields a low win rate (32.1%) and negative CTR improvement (-9.83%), revealing the gap between model generation quality and online deployment effectiveness. Top-1 sampling improves win rate through only the diverse ad sampling strategy, but still results in a slight CTR reduction (-0.6%). This suggests that direct imitation does not ensure suitability for the target item, underscoring the need for preference optimization from online feedback.
In contrast, preference alignment methods leverage online feedback for model optimization, all achieving higher win rates ($>$50%) over the human-crafted baseline. Among these methods, DPO outperforms Top-1 SFT, demonstrating the advantage of reinforcement learning to align with user engagement. Our method CTOP extends DPO by introducing dynamic weights of preferences based on CTR gain and confidence, achieving superior performance. Through prioritizing high-confidence, high-gain pairs, the model learns from the most meaningful feedback, leading to more robust and effective ad text generation.
@@
We conduct an ablation to evaluate the impact of gain and confidence coefficients in our CTR-driven preference optimization. As shown in Table 5, removing both coefficients reduces the method to standard DPO, which yields a 56.0% win rate and +3.48% relative CTR improvement. Adding only the confidence coefficient improves the win rate to 58.5% and CTR to +3.79%, indicating better stability by prioritizing reliable preference pairs. Our full method, which incorporates both gain and confidence, further boosts the win rate to 60.2% and relative CTR improvement to +4.76%. These results demonstrate that jointly considering both factors produces the best overall performance: gain emphasizes the impact of preference pairs on CTR, while confidence mitigates noise from unreliable data.
@@
We deployed our proposed CTOP for item title generation on a large-scale e-shopping platform across various ad scenarios. In initial A/B/n testing, the candidate group showed a 0.46% CTR decrease versus the human-crafted baseline, indicating the necessity of preference alignment. After CTR-driven optimization, the generated titles achieved a relative improvement of +1.11% in CTR and +1.02% in RPM (revenue per mille) over the human-crafted baseline under full platform traffic, since late 2024. The significant performance improvement demonstrates the strong applicability and effectiveness of CTOP in real-world advertising systems.
@@
Figure 6 provides comparison cases between human-crafted item titles and our generated titles on the shopping platform. The human-crafted titles are often lengthy, keyword-stuffed, lacking readability or user appeal. In contrast, our generated titles are concise and stylistically diverse—ranging from benefit-driven to persona-targeted—using engaging language that is well-aligned with both item features and user attention, resulting in better CTR performance.
@@ Conclusion
In this work, we propose CTOP, a novel CTR-driven ad text generation method via online feedback preference optimization. CTOP employs a two-stage framework, comprising a diverse ad text sampling and a CTR-driven preference optimization approach, tailored for the low-quality sampling-based decoding and noisy online feedback, respectively. Extensive experiments have demonstrated the effectiveness of our method in both offline and online metrics, significantly improving CTR in real-world advertising scenarios. This study contributes an effective solution to the industry and inspires further research in e-commerce advertising.

### v3

@@ Introduction
Advertising text is crucial in online advertising, appearing in almost every ad. Effective ad texts not only feature well-crafted expression and accurate conveyance of item information (Murakami et al. 2022), but also achieve strong performance in terms of click-through rate (CTR). Large Language Models (LLMs) enable automated ad text generation (ATG) (Mita et al. 2024) from basic item information and simple instructions, offering significantly higher efficiency than manual creation. Some works (Golobokov et al. 2022; Chai et al. 2022; Jin et al. 2023; Kamigaito et al. 2024) focus on offline metrics like diversity or fluency to enhance the quality of generated ad texts. However, these methods do not guarantee higher CTR performance compared to human-crafted texts, revealing a gap between offline metrics optimization and online deployment effectiveness.
@@
Figure 1: Challenges in high-CTR ad text generation. (a) Poor diversity and inferior CTR performance of sampling-based decoding strategy. (b) Noisy online feedback due to uneven page view ($PV$) distribution across items. (c) We propose an innovative two-stage framework: (1) diverse ad text sampling and (2) CTR-driven preference optimization tailored for these challenges.
@@
Recently, studies (Kanungo et al. 2022; Wei et al. 2022; Murakami et al. 2025) have focused on directly optimizing for online metrics, particularly CTR. CTR has long been regarded as a key metric for evaluating online advertising effectiveness, as clicks serve as the prerequisite for subsequent conversion to other revenue-related metrics. These approaches typically generate multiple ad texts via sampling-based decoding from LLMs for online testing. The generation models are subsequently optimized using online feedback through algorithms such as contrastive learning, semi-supervised learning, and reinforcement learning.
@@
However, these approaches often suffer from the low quality of sampling-based decoding and the noise in online feedback, making it difficult to achieve a significant CTR advantage over human-crafted ad texts. We conduct some analytical experiments and reveal the challenges to realize the CTR improvement. Figure 1(a) shows that generating multiple ad texts via sampling-based decoding from LLMs results in poor diversity and inferior CTR performance. Figure 1(b) illustrates data confidence issues arising from the uneven distribution of item page views ($PV$) in online feedback. Short-term online testing makes it difficult to accumulate sufficient high-PV data, causing substantial noise in online feedback. These challenges constrain further improvements in CTR optimization of ATG tasks.
@@
In this work, we propose CTOP, a novel CTR-driven ad Text generation method via Online feedback Preference optimization. CTOP employs an innovative two-stage framework consisting of diverse ad text sampling and CTR-driven preference optimization for high-CTR ad text generation, as illustrated in Figure 1(c). In the first stage, we design a diverse ad text sampling strategy to address the low quality of sampling-based decoding. Specifically, we generate candidate ad texts via a style transfer task, imitating the style of high-quality exemplars. For each item, the strategy uses retrieval-augmented generation (RAG) to retrieve multiple chain-of-thought (CoT)-augmented ad text exemplars, each serving as a one-shot in-context example. In the second stage, we conduct CTR-driven preference optimization tailored for noisy online feedback. We select winning candidates that outperform their corresponding human-crafted ad texts in online A/B/n testing and construct preference pairs by treating the former as preferred and the latter as dispreferred. Each pair is assigned a training weight calculated from gain and confidence coefficients based on its CTR improvement and reliability.
@@
Extensive experimental results on offline and online metrics demonstrate the effectiveness of CTOP. Our method enables efficient candidate generation of diverse high-quality ad texts and better aligns with online user feedback during preference optimization. Notably, we have successfully applied our method to the item title generation task on a large-scale online shopping platform across various advertising display scenarios. Since late 2024, our approach has achieved a relative increase of 1.11% in click-through rate (CTR) and 1.02% in revenue per mille (RPM) compared to the human-crafted baseline, showcasing its strong applicability and effectiveness in advertising systems. The main contributions of our work can be summarized as follows:
@@
  • We propose a novel ad text generation method which optimizes for CTR via preference optimization from online feedback, closing the gap between generation quality and online performance of ad texts.
@@
  • We design an innovative two-stage framework, consisting of (1) diverse ad text sampling via one-shot in-context learning, using RAG to provide high-quality exemplars with CoTs; and (2) CTR-driven preference optimization method that dynamically weighs training pairs by gain and confidence coefficients based on online feedback.
@@
However, evaluating whether a generated ad text achieves “high CTR” is challenging without a meaningful baseline. In practice, each item has a human-crafted ad text $y_{\text{hu}}$ typically written by merchants or the advertising platforms based on domain expertise or empirical knowledge. These human-crafted ad texts generally achieve a certain level of quality and CTR, which serve as a strong and interpretable baseline. Therefore, any generated ad text $y$ that outperforms $y_{\text{hu}}$ in online testing can be considered achieving high CTR. To this end, we reformulate the objective as maximizing the CTR improvement over human-crafted ad texts:
@@
Optimizing ad text generation for high CTR aligns naturally with reinforcement learning (RL), where the goal is to maximize user engagement in real-world environments. In our setting, online A/B/n testing yields an off-policy dataset comprising both LLM-generated and human-crafted texts. Moreover, CTR feedback is delayed, noisy, and requires significant online accumulation, making it challenging to train an explicit reward model for CTR prediction or preference estimation. These factors motivate us to adopt off-policy RL methods with implicit reward modeling (Rafailov et al. 2023; Pal et al. 2024; Wu et al. 2024; Azar et al. 2024; Kim et al. 2024; Ethayarajh et al. 2024; Liu et al. 2025).
Here we adopt DPO (Rafailov et al. 2023) due to its simplicity and efficiency. DPO operates on preference pairs $(x,y_{w},y_{l})$, where $y_{w}$ and $y_{l}$ are the preferred and dispreferred responses with input context $x$. Given a policy model $\pi_{\theta}$ and a reference model $\pi_{\mathrm{ref}}$, the DPO objective is:
@@
We construct preference pairs with each winning candidate as $y_{w}$ and its human-crafted ad text as $y_{l}$ from online feedback. A Qwen-2.5-7B model is fine-tuned on diverse candidate ad texts as $\pi_{\mathrm{ref}}$. With a dataset of 50K pairs from online A/B/n testing, we train a DPO-optimized model. However, A/B testing against human-crafted ad texts yields only a 42.1% win rate.
@@
In this section, we introduce CTOP, a novel CTR-driven ad Text generation method via Online feedback Preference optimization. To ensure meaningful improvements in online performance, we formulate the task objective as maximizing the CTR improvement over human-crafted ad texts, defined in Equation 2. As illustrated in Figure 4, our method employs an innovative two-stage framework of diverse ad text sampling and CTR-driven preference optimization tailored for the two challenges. To address the low-quality issue of sampling-based decoding, we design a diverse ad text sampling strategy in the first stage. To better learn from noisy online feedback, we perform CTR-driven preference optimization in the second stage. The following subsections elaborate on each stage of our proposed method.
@@
As depicted in Challenge I, candidate ad text generation via sampling-based decoding causes poor diversity and inferior CTR performance. To effectively generate candidates of both high quality and true diversity, we design a diverse ad text sampling strategy, shown in Figure 4(a). The core of our strategy is to formulate candidate generation as a style transfer task, where the model imitates high-quality exemplars via one-shot in-context learning.
To support this, we first construct a repository of high-quality ad text exemplars. These exemplars are either crafted by domain experts or selected from successful online cases with proven strong CTR performance. Each exemplar consists of the item information, the ad text, and a chain-of-thought (CoT) with imitation tips to guide faithful style transfer. We pre-generate offline CoTs for all exemplars in the repository for efficiency.
@@
Owing to the inherent stylistic diversity among high-quality exemplars in the repository, our strategy is able to maintain both quality and diversity even as the number of candidates per item $K$ increases. Moreover, this diversity is controllable: our strategy can generate new styles of ad texts simply by collecting additional exemplars without extra training. Compared to methods that demand manual prompt engineering for each new style, our strategy offers a more scalable and efficient solution with minimal overhead.
@@
Following our diverse ad text sampling strategy, we deploy the generated candidates and the human-crafted ad texts for short-term online A/B/n testing, as shown in Figure 4(b). A candidate that achieves a higher CTR than its corresponding human-crafted ad text is defined as a winning candidate. We construct preference pairs with each winning candidate as $y_{w}$ and its human-crafted ad text as $y_{l}$. However, our analysis in Challenge II reveals that noisy online feedback negatively impacts the quality of preference optimization. To address this issue, we propose a CTR-driven preference optimization approach illustrated in Figure 4(c), where each pair is assigned a dynamic weight calculated from gain and confidence coefficients.
@@
We conduct online experiments on a large-scale e-shopping platform, with the task of generating item titles across display advertising scenarios. For diverse ad text sampling, we first build a high-quality item title repository with offline CoTs. A Qwen-2.5-7B model is trained on the ad text style transfer task as the style transfer model. Then we retrieve the top-5 relevant exemplars per item and use each as a one-shot in-context example for style transfer. We select 200K items across different categories for diverse candidate generation, resulting in a total of 1 million candidate ad titles. These candidates, together with the human-crafted item titles, are deployed for a two-week online A/B/n testing using 5% of platform traffic as the experimental group. A parallel AA group is simultaneously deployed with identical content and traffic allocation.
For CTR-driven preference optimization, we first construct a preference dataset of 500K pairs by pairing each winning candidate and its corresponding human-crafted title from the online A/B/n testing. Each pair is assigned a dynamic weight based on its gain coefficient (Equation 5, with scaling factor $m=2.5$) and confidence coefficient (Equation 6). We use a Qwen-2.5-7B model fine-tuned on diverse candidate titles as the reference model $\pi_{\text{ref}}$, which also initializes the policy model $\pi_{\theta}$. We optimize the objective in Equation 7 for one training epoch with a learning rate of $2\times 10^{-5}$. All training processes are conducted on machines equipped with 8 NVIDIA A100 GPUs.
Due to business sensitivity concerns, we are unable to publicly release the data used in the experiments. See Appendix for more implementation details.
@@
  • Top-1 sampling: Our diverse ad text sampling with sampled number $K=1$ (i.e., the most related exemplar).
@@
w/o preference alignment
Prompt-only 41.0 -9.83
@@
For baselines without preference alignment (Prompt-only and Top-1 sampling), both underperform the human-crafted ad texts. Prompt-only yields a low win rate (41.0%) and negative CTR improvement (-9.83%), revealing the gap between generation quality and online performance of ad texts. Top-1 sampling improves win rate through only the diverse ad text sampling strategy, but still results in a slight CTR reduction (-0.6%). This suggests that style transfer does not ensure suitability for the target item, underscoring the need for preference optimization from online feedback.
In contrast, preference alignment methods leverage online feedback for model optimization, all achieving higher win rates ($>$50%) over the human-crafted baseline. Among these methods, DPO outperforms Top-1 SFT, demonstrating the advantage of reinforcement learning to align with user engagement. Our method CTOP extends DPO by introducing dynamic weights of preference pairs based on CTR gain and confidence, achieving superior performance. Through prioritizing high-confidence, high-gain pairs, the model learns from the most meaningful feedback, leading to more robust and effective ad text generation.
@@
We conduct an ablation to evaluate the impact of gain and confidence coefficients in our CTR-driven preference optimization. As shown in Table 5, removing both coefficients reduces the method to standard DPO, which yields a 56.0% win rate and +3.48% relative CTR improvement. Adding only the confidence coefficient improves the win rate to 58.5% and CTR to +3.79%, indicating better stability by prioritizing reliable preference pairs. Our full method, which incorporates both gain and confidence, further boosts the win rate to 60.2% and relative CTR improvement to +4.76%. These results demonstrate that jointly considering both factors produces the best overall performance: gain emphasizes the impact of preference pairs on CTR improvement, while confidence mitigates noise from unreliable data.
@@
We deployed our proposed CTOP for item title generation on a large-scale e-shopping platform across various ad scenarios. In initial A/B/n testing, the candidate ad texts showed a -0.46% CTR decrease versus the human-crafted baseline, indicating the necessity of preference alignment. After CTR-driven optimization, the generated titles achieved a relative improvement of +1.11% in CTR and +1.02% in RPM (revenue per mille) over the human-crafted baseline under full platform traffic, since late 2024. The significant performance improvement demonstrates the strong applicability and effectiveness of CTOP in real-world advertising systems.
@@
Figure 6 provides comparison cases between human-crafted item titles and our generated titles on the shopping platform. Human-crafted titles are often informative but keyword-stuffed, lacking readability and user appeal. In contrast, our generated titles are concise and stylistically diverse—ranging from benefit-driven to persona-targeted—using engaging language well-aligned with both item features and user attention, resulting in better CTR.
@@ Conclusion
In this work, we propose CTOP, a novel CTR-driven ad text generation method via online feedback preference optimization. CTOP employs a two-stage framework, comprising diverse ad text sampling strategy and CTR-driven preference optimization, tailored for the low-quality sampling-based decoding and noisy online feedback, respectively. Extensive experiments have demonstrated the effectiveness of our method on both offline and online metrics, significantly improving CTR in real-world advertising scenarios. This study contributes an effective solution to the industry and inspires further research in e-commerce advertising.
Supplementary Material
Appendix A Appendices
This supplementary material provides:
  1. 1.
    Exemplars from the High-Quality Ad Text Repository: Illustrative cases showing several exemplars from the high-quality ad text repository in our diverse ad text sampling.
  2. 2.
    Details of Training Tasks: Implementation details of the training tasks in our method, including the style transfer model, the reference model, and the policy model .
  3. 3.
    Additional Case Studies: More detailed comparisons among human-crafted ad texts, underperforming candidates in diverse ad text sampling, and final outputs after our CTR-driven preference optimization, demonstrating the effectiveness of our proposed method.
Appendix B Exemplars from the High-Quality Ad Text Repository
In our diverse ad text sampling strategy, the high-quality ad text repository serves as a core component. This repository consists of high-quality ad texts either crafted by domain experts or selected from successful online cases, with strong CTR performance proven beforehand. Table 6 presents several exemplars from the repository. Each exemplar in the repository details the item, the corresponding ad text, and the chain-of-thought (CoT) serving as imitation tips for style transfer. The imitation tips provide a structural decomposition and a step-by-step writing guidance to capture the stylistic essence of the exemplar. To address the inefficiency of online CoT during inference, we pre-generate offline CoTs for all exemplars in the repository. Specifically, we use GPT-4 for CoT generation with the Offline CoT Generation prompt template (Table 8). By introducing high-quality exemplars and their imitation tips as one-shot in-context examples, our strategy enables controllable style transfer and improves the diversity and effectiveness of candidate ad text generation.
Appendix C Details of Training Tasks
Our method involves 3 models to be trained: (1) the style transfer model in diverse ad text sampling, (2) the reference model and (3) the policy model in CTR-driven preference optimization. Table 6 presents the training details for each model used in our method, including the base model, the format, and the size of the training data. Due to the business information involved in the training data, we are unable to publicly release the data used in our experiments.
The style transfer model is trained via supervised fine-tuning on instruction-response pairs. The instruction is constructed by filling the Style Transfer prompt template (Table 8) with an exemplar ad text <exemplar> from the repository, its corresponding CoT <imitation tips>, and the target item information <item info>. The response is the generated ad text that imitates the style of the exemplar for the target item. To construct the training data, we select 1,000 items disjoint from those in the high-quality ad text repository. For each item, we randomly sample 10 exemplars and use GPT-4 to generate style-imitated ad texts using the same prompt template, resulting in 10K instruction-response pairs. A Qwen-2.5-7B model is fine-tuned on this dataset for one epoch with a learning rate of $2\times 10^{-5}$, yielding the final style transfer model.
The reference model is trained via supervised fine-tuning on instruction-response pairs to serve as a behavioral prior for subsequent preference optimization. The instruction is constructed by filling the Ad Text Generation prompt template (Table 8) with the target item information <item info>. The response is the generated ad text for the target item. We select 100K generated candidates in the diverse ad text sampling to construct the training data. A Qwen-2.5-7B model is fine-tuned on this dataset for one epoch with a learning rate of $2\times 10^{-5}$, yielding the reference model.
The policy model, initialized from the reference model, is trained on preference pairs consisting of instruction $x$, preferred response $y_{w}$, and dispreferred response $y_{l}$. The instruction is constructed by the same Ad Text Generation prompt template with the target item information <item info>. We construct 500K preference pairs from the online A/B/n testing feedback, with each winning candidate of an item as $y_{w}$ and its human-crafted ad text as $y_{l}$. Then, we leverage our CTR-driven preference optimization (Equation 7) to train the policy model for one epoch, with a learning rate of $2\times 10^{-5}$ and scaling factor $m=2.5$ (Equation 5).
Appendix D Additional Case Studies
Figure 7 presents detailed comparisons among human-crafted ad texts, underperforming candidates from diverse ad text sampling, and final outputs after our CTR-driven preference optimization. Human-crafted texts are informative but often lack persuasive elements such as emotional hooks. Diverse ad text sampling improves variation but may generate inappropriate candidates that fail to align with user preferences. In contrast, our method generates more attractive and targeted ad texts by learning high-CTR patterns from online feedback. For instance, it better utilizes bracketed emphasis (e.g., [IPX7 Waterproof]) to highlight key selling points. Moreover, our model learns a better keyword-placing order to place attention-grabbing keywords or essential features earlier (e.g., [304 Stainless] in the beginning). These cases demonstrate that our method generates ad texts that are well-aligned with both item features and user preferences, resulting in better CTR performance.
Item Exemplar Ad Text Imitation Tips (Chain-of-Thought)
Back Support Belt Say Goodbye to Back Pain Start with This Belt All-Day Comfort and Reliable Support #Structure targeted problem + action guidance + item name + additional information #Writing Guidance Start with the specific problem or user pain point being addressed. Follow with an action-oriented guidance phrase. State the item name. Add supplementary information to highlight key features or advantages.
Shirt Dress Comfy|Breathable|Slimming-Vintage Print Shirt Dress #Structure benefit1|2|3-item description #Writing Guidance Start with top 3 benefits using short keywords. List benefits in descending order of importance. Separate keywords with vertical bars "|". Item description includes the item name and other key attributes. Use a dash "-" to connect the final benefit to the item description.
Women’s Heel [Petite Height Boost] Women’s Chunky Heel, Waterproof Platform, French Design #Structure [core benefit] + item description #Writing Guidance Start with the main selling point or targeted benefit, enclosed in brackets "[]". Item description includes the item name and other key attributes. Organize keywords of item description in descending order of importance.
Table 6: Illustrative exemplars from the high-quality ad text repository. Each exemplar lists the item, the ad text, and the imitation tips (CoT) for text style transfer, including a structural decomposition and a step-by-step writing guidance.
Model Base Model Data Format Template Data Size
Style Transfer Model Qwen-2.5-7B instruction-response [Style Transfer] 10K
Reference Model Qwen-2.5-7B instruction-response [Ad Text Generation] 100K
Policy Model Reference Model preference pairs $(x,y_{w},y_{l})$ [Ad Text Generation] 500K
Table 7: Training details for each model used in our framework. The table summarizes for each model: the base model, data format used for training, the corresponding prompt template (see Table 8), and the data size.
Task Prompt Template
Offline CoT Generation You are an experienced e-commerce ad specialist. Given a exemplar ad text and its corresponding item information, please analyze the ad text and write some imitation tips for ad text style transfer. # Exemplar Ad Text <exemplar> # Item Information <item info> #Rules - The imitation tips consist of a structural decomposition of the given ad text and a writing guidance for style transfer. - Structural Decomposition: Provide an equation-like summary that reflects the overall structure of the ad text (e.g., benefit 1 + benefit 2 + item name + key attributes). - Writing Guidance: Present, as an unordered list, actionable steps or suggestions on how to reproduce this style when generating new ad texts for other items. - Output the result in the following format: # Structure <structural decomposition> # Writing Guidance <unordered guidance list>
Style Transfer You are an experienced e-commerce ad specialist. Please generate a new ad text for the target item by imitating the style of the given exemplar. # Exemplar ad text <exemplar> # Style analysis <imitation tips> # Target item information <item info> # Notice - Do not fabricate any details (e.g., price, promotion, shipping, endorsements) not present in item information. - Output only the generated ad text without extra explanation.
Ad Text Generation You are an experienced e-commerce ad specialist capable of creating attractive ad texts that significantly enhance user interest and click-through rate. Please generate an ad text for the target item. # Target item information <item info> # Notice - Do not fabricate any details (e.g., price, promotion, shipping, endorsements) not present in item information. - Output only the generated ad text without extra explanation.
Table 8: Prompt templates for the offline CoT generation, the style transfer, and the ad text generation task.
Figure 7: Detailed comparisons among human-crafted ad texts, underperforming candidates in diverse ad text sampling, and final outputs after our CTR-driven preference optimization.